# Semialgebraic Neural Networks: From roots to representations

### v1

@@ A.1 Inf-sup definable piecewise polynomials and the Pierce–Birkhoff conjecture
def-piecewise-polynomial Let $k,n,J\in\mathbb{N}$, and $D\subset\mathbb{R}^{n}$. Then $f:D\to\mathbb{R}$ is *piecewise polynomial*, or $f\in PWP(D)$, if there exists a collection of semialgebraic sets $\{S_{i}\}_{i=1}^{J}$ such that $D\subset\bigcup_{i=1}^{J}S_{i}$, and a collection of polynomials $\{p_{i}\}_{i=1}^{J}\subset\mathbb{R}[X_{1},\dots,X_{n}]$ such that $f|_{S_{i}}=p_{i}$ for all $i=1,\dots,J$.
@@
The case $n=2$ was proved by Jacek Bochnak and Gustave Efroymson in [Bochnak et al., 1998], and then again by Louis Mahé [Mahé, 1984]. Mahé has also shown that the conjecture holds when $n=3$ up to arbitrarily small neighborhoods of finitely many points. See Madden [2011] for further discussion of the Pierce–Birkhoff conjecture and related problems.
Later, we will show that MRNNs can exactly compute any continuous piecewise polynomial if and only if the Pierce–Birkhoff conjecture holds.
@@ Appendix B Matrix-recurrent neural networks
Semialgebraic Neural Networks (SANNs) are built on top of a related architecture called Matrix-Recurrent Neural Networks (MRNNs), a type of polynomial network with ReLU activations. This chapter introduces MRNNs and exactly characterize their range as inf-sup definable (ISD) piecewise polynomials. Our analysis of the expressivity of MRNNs are inspired by Balestriero and Baraniuk [2021].
@@ B.1 Architecture
In this work, we study a type of Operator Recurrent Neural Network (ORNN) [de Hoop et al., 2022] where the input to the network is a matrix $X\in\mathbb{R}^{n\times m}$ and initial vector $x\in\mathbb{R}^{k_{\text{in}}}$, and the output is a vector $f(X,x)\in\mathbb{R}^{k_{\text{out}}}$. To emphasize that our results hold specifically for matrix inputs rather than general operators, we call this subset of ORNN’s the Matrix Recurrent Neural Networks (MRNNs). The architecture of an $L$-layer MRNN is defined
@@
where $(I\otimes X)$ denotes a Kronecker product with an identity matrix $I$ whose size may vary in each layer. Notice in particular that the input matrix $X$ is inserted multiplicatively into each layer. Our definition of an MRNN matches the original definition of a width-expanded ORNN with matrix input presented in de Hoop et al. [2022]. As noted in that paper, the restriction on $h_{0}$ does not affect the range of $f$ since $h_{1}$ will invariably equal $b_{1,0}$, which is learned.
@@
def-MRNN Let $k,L\in\mathbb{N}$, and $D$ be a subset of $\mathbb{R}^{m\times n}$. Then $f:D\to\mathbb{R}^{k}$ is in $MRNN_{L}(D,\mathbb{R}^{k})$ if $f$ can be computed via equations (23-25).
@@
In contrast to standard ReLU networks, we will show that MRNNs for a ring of piecewise polynomial functions. The kernels of these functions—that is, the set of points $X$ such that $f(X)=0$—form semialgebraic sets in $\mathbb{R}^{m\times n}$.
@@ B.2 Range
Numerous authors have observed that linear ReLU networks compute multivariate linear splines, where each linear region is a polyhedron defined by the combination of active neurons, that is, where $b_{\ell}+A_{\ell}h_{\ell-1}\geq 0$—see, for example, [Montúfar et al., 2014, Wang et al., 2019]. Likewise, de Hoop et al. [2022] showed that MRNNs (or rather ORNNS in general) compute piecewise polynomials, where the semialgebraic decomposition of the domain is similarly defined by the combination of active neurons
$b_{\ell,1}+A_{\ell,1}h_{\ell-1}+B_{\ell,1}(I\otimes X)h_{\ell-1}$
@@
By definition of $ISD(\mathbb{R}^{mn}),$ $f$ is a lattice polynomial $q$ of some $\{p_{1},\dots,p_{J}\}\subset\mathbb{R}[X_{1},\dots,X_{mn}]$. An easy construction shows there exists networks $\mathcal{N}_{1},\dots,\mathcal{N}_{J}$ computing the polynomials $\mathcal{N}_{j}=p_{j}\circ\mathrm{vec}^{-1}$. Another easy construction shows that $MRNN(\mathbb{R}^{m\times n})$ is closed under the lattice operations, so the lattice polynomial $q$ can be computed by a network. In particular, $\mathcal{N}=q\left(\mathcal{N}_{1},\dots,\mathcal{N}_{J})\right)$ is the desired network. ∎
@@
$h_{L}=b_{L,0}+A_{L,0}h_{L-1}+B_{L,0}(I\otimes X)h_{L-1}+\sigma\left(b_{L,1}+A_{L,1}h_{L-1}+B_{L,1}(I\otimes X)h_{L-1}\right).$ (26)
@@
for some $N\in\mathbb{N}$ corresponding to the width of this layer. For every $j=1,\dots,N$, $g_{j}(X)$ is computed by an $L-1$ network, so the inductive hypothesis states there exists $f_{j}\in ISD(\mathbb{R}^{mn})$ such that $g_{j}=f_{j}\circ\mathrm{vec}$. We now rewrite each term of equation equation 26 using these $f_{j}$. Let $i\in\{0,1\}$.
@@
We can thus rewrite equation 26 in the equivalent form
@@
Theorems LABEL:thm-MRNN-equals-ISD and LABEL:thm-lattice-over-subring-of-f-ring-is-ring allow us to conclude that $MRNN(\mathbb{R}^{n\times m},\mathbb{R}^{k})$ forms a unital, associative algebra over $\mathbb{R}$; that is, every scalar multiple, sum, and Hadamard product of a MRNNs is representable by an MRNN. Associativity and existence of a multiplicative identity, as well as the ability to compute scalar multiples and sums are obvious and hold for linear ReLU networks as well. What distinguishes MRNNs from linear ReLU networks is the ability to compute Hadamard products. We require this property to build SANNs capable of computing any semialgebraic function in the next chapter.
While the proofs above are not explicitly constructive, they can be made constructive using the identities in Henriksen and Isbell [1962] or Hager and Johnson [2010]. Unfortunately, representing the Hadamard product of two MRNNs as an MRNN requires exponentially many parameters, which limits the practical usefulness of a direct application of these constructions in code.
In their original paper introducing the architecture, de Hoop et al. [2022] note that the range of an MRNN is a multivariate piecewise polynomial; we have refined this result to show that MRNNs are ISD. We now ask the natural question, is continuous every piecewise polynomial representable by an MRNN? As noted previously, this question is the famous Pierce–Birkhoff conjecture, a long-standing open problem in real algebraic geometry. MRNNs can represent every piecewise polynomials if and only if the Pierce–Birkhoff conjecture is true.
@@ Appendix C Semialgebraic Functions as Kernels of ISD Functions
Theorem LABEL:prop-semialgebraic-sets-as-kernels.
@@
Each term $p_{i}^{2}(x)$ (resp. $\max\{0,-q_{i}(x)\}^{2}$) is zero precisely when the corresponding condition $p_{i}(x)=0$ (resp. $q_{i}(x)\geq 0$) is satisfied. Also, the function $q\to\max\{0,q\}^{2}$ is in $C^{1}$ and has the continuous derivative $q\to 2\max\{0,q\}$. Moreover, each term is non-negative, so the sum is non-negative and zero if and only if every term is zero, meaning every condition is satisfied simultaneously. Finally, $f$ is ISD since it is constructed using the ISD lattice-ring operations on polynomials (see Appendix A), and it is $C^{1}$ since each term in the sum is $C^{1}$.
Now let $S$ be an arbitrary closed semialgebraic set; i.e. the union of basic semialgebraic sets. We may take the basic sets to be closed, since
@@
Theorem LABEL:cor-semialgebraic-graphs-as-kernels.
@@ D.1 Semialgebraic Sard’s and Transversality Theorems
In the “textbook” Sard’s theorem, the smoothness requirements on the function depend on the dimension of the range [Abraham and Robbin, 1967]. In contrast, for semialgebraic functions continuous differentiability is sufficient for the theorem to hold. A proof can be found in Kurdyka et al. [2000].
@@
A proof can be found in [Abraham and Robbin, 1967], substituting the semialgebraic Sard’s theorem for Smale’s theorem as required.
@@ D.2 Homotopy existence
We now state a standard result in the theory of homotopy continuation for solving nonlinear equations. Theorem LABEL:thm-homotopy-continuation is adapted from Krantz and Parks [2003] with small modifications using the theorems from Appendix D.1 to specialize to $ISD^{1}$ and avoid any $C^{\infty}$ assumptions. It provides sufficient conditions under which we are guaranteed to be able to evaluate a homotopy from $t=0$ until $t=1$.
@@ D.3 Curve tracing
Let $z(s):=\big{(}y(s),t(s)\big{)}$, and $H\in ISD^{1}(\mathbb{R}^{n}\times\mathbb{R},\mathbb{R}^{n})$ satisfy the hypotheses of Theorem LABEL:thm-homotopy-continuation. Then $(y,1)\in\ker H$ can be computed by solving the ODE initial-value problem defined by the “arc-length parameterization” [Chen and Li, 2015]:
@@
With proper choice of $\mathcal{N}$ and $c_{\text{max}}$, the curve-tracing ODE system (31)-(34) can be written in the form of the ODE system (7)-(8) that define SANNs. At time $s=0$, write
@@
where $\alpha$ is a constant chosen to satisfy (32) and (33). The $M$ and $b$ defined above appear in the ODE (7)-(8. Thus we have shown for $H$ satisfying the hypotheses of Theorem LABEL:thm-homotopy-continuation, $y_{0}=0$, and large enough $c_{\max},$ there is an $\mathcal{N}\in ISDnet(m,n,k)$ such that the integral curve traced in solving the ODE (7)-(8) is “correct” at time $s=0$; i.e. it is tangent to the integral curve defined by $\dot{z}$ solving the IVP (31)–(34).
For time $s>0$, we are no longer guaranteed that the first $n$ columns of $DH$ form an invertible submatrix. However, by assumption A.3, $DH$ always has rank $n$, so every $s$ belongs to some interval where we can choose a column of $DH$ to be $\widehat{b}$, and the remaining columns to be $\widehat{M}$. The computations in each of these intervals can be glued together in a way consistent with the ODE (7)-(8) used to compute the entire integral curve. The details can be found in the proof of Theorem 12 in Appendix E.
@@ E.1 Notation
The theorems in this section require us to show the existence of ISD $\mathcal{N}$ generating certain initial value problems for (7)-(8). In Algorithm 1, $\dot{z}$ closes over the inputs $\mathcal{N}$, $c_{\max}$, and $x$, and ODESolve always runs from time $s=0$ to $s=1$. For this section, we instead make the dependence on all variables explicit by using the signature
@@ E.2 Glueing
For $t$-glueing, we build $\mathcal{N}_{3}$ to interpolate between $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ in disjoint closed $t$ intervals $[t_{0},t_{1}]$ and $[t_{2},t_{3}]$. We use $t$-glueing when we can guarantee the integral curve traced by ODESolvewill never enter the region $(t_{1},t_{2})$, so the value of $\mathcal{N}_{3}$ may be arbitrary here. We are free to use it to interpolate between $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$.
@@
where $u$ is the Hermite interpolant [Süli and Mayers, 2003] such that
@@
Simple calculus verifies conclusion 37, and conclusion 38 is clear. ∎
@@
Choose $\epsilon>0$ such that $s_{0}<s_{1}-\epsilon<s_{1}+\epsilon<s_{\text{final}}$, and apply 38 to $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ so that both $b_{1}$ and $b_{2}$ are identically 00 in the region $(s_{1}-\epsilon,s_{1}+\epsilon).$ No matter the value of $M_{1}$ and $M_{2}$, both $\dot{y}$ and $\dot{t}$ are identically 00. Let $\lambda(s)=(s-(s_{1}-\epsilon))/(2\epsilon)$. Define
@@
$M_{3}$ is clearly continuous, and $b_{3}$ is continuous since $b_{1}$ and $b_{2}$ are both 00 when $s=s_{1}.$ The constructed $\mathcal{N}_{3}$ is ISD and satisfies conclusion equation 45. ∎
@@
Given $\mathcal{N}_{1}$, $\mathcal{N}_{2}$, $c_{1}$, $c_{2}$, $y_{1}$, $y_{2}$, $t_{1}$, $t_{2}$,$s_{0}$, $s_{\text{final}}$, and let
@@
Suppose $\mathcal{T}_{1}$ are disjoint $\mathcal{T}_{2}$. Then there exists $\mathcal{N}_{3}$, $c_{3}$ such that

### v2

@@ A.1 Inf-sup definable piecewise polynomials and the Pierce–Birkhoff conjecture
def-piecewise-polynomial Let $k,n,J\in\mathbb{N}$, and $D\subset\mathbb{R}^{n}$. Then $f:D\to\mathbb{R}$ is piecewise polynomial, or $f\in PWP(D)$, if there exists a collection of semialgebraic sets $\{S_{i}\}_{i=1}^{J}$ such that $D\subset\bigcup_{i=1}^{J}S_{i}$, and a collection of polynomials $\{p_{i}\}_{i=1}^{J}\subset\mathbb{R}[X_{1},\dots,X_{n}]$ such that $f|_{S_{i}}=p_{i}$ for all $i=1,\dots,J$.
@@
The case $n=2$ was proved by Jacek Bochnak and Gustave Efroymson (Bochnak et al., 1998), and then again by Louis Mahé (Mahé, 1984). Mahé has also shown that the conjecture holds when $n=3$ up to arbitrarily small neighborhoods of finitely many points. See Madden (2011) for further discussion of the Pierce–Birkhoff conjecture and related problems.
Later, we will show that the neural network architecture introduced in the next section can exactly compute any continuous piecewise polynomial if and only if the Pierce–Birkhoff conjecture holds.
@@ Appendix B Matrix-recurrent neural networks
Semialgebraic Neural Networks (SANNs) are built on top of a related architecture called Matrix-Recurrent Neural Networks (MRNNs), a type of polynomial network with ReLU activations. This chapter introduces MRNNs and exactly characterize their range as inf-sup definable (ISD) piecewise polynomials. Our analysis of the expressivity of MRNNs is inspired by Balestriero & Baraniuk (2021).
@@ B.1 Architecture
In this work, we study a type of Operator Recurrent Neural Network (ORNN) (de Hoop et al., 2022) where the input to the network is a matrix $X\in\mathbb{R}^{n\times m}$ and the output is a vector $f(X)\in\mathbb{R}^{k_{\text{out}}}$. To emphasize that our results hold specifically for matrix inputs rather than general operators, we call this subset of ORNN’s the Matrix Recurrent Neural Networks (MRNNs). The architecture of an $L$-layer MRNN is defined
@@
where $(I\otimes X)$ denotes a Kronecker product with an identity matrix $I$ whose size may vary in each layer. Notice in particular that the input matrix $X$ is inserted multiplicatively into each layer. Our definition of an MRNN matches the original definition of a width-expanded ORNN with matrix input presented in de Hoop et al. (2022). As noted in that paper, the restriction on $h_{0}$ does not affect the range of $f$ since $h_{1}$ will invariably equal $b_{1,0}$, which is learned.
@@
def-MRNN Let $k,L\in\mathbb{N}$, and $D$ be a subset of $\mathbb{R}^{m\times n}$. Then $f:D\to\mathbb{R}^{k}$ is in $MRNN_{L}(D,\mathbb{R}^{k})$ if $f$ can be computed via equations (22-24).
@@
In contrast to standard ReLU networks, we will show that MRNNs form a ring of piecewise polynomial functions. The kernels of these functions—that is, the set of points $X$ such that $f(X)$ vanishes—form semialgebraic sets in $\mathbb{R}^{m\times n}$.
@@ B.2 Range
Numerous authors have observed that linear ReLU networks compute multivariate linear splines, where each linear region is a polyhedron defined by the combination of active neurons—that is, where $b_{\ell}+A_{\ell}h_{\ell-1}\geq 0$ (see, for example, Montúfar et al. (2014) and Wang et al. (2019)). Likewise, de Hoop et al. (2022) showed that MRNNs (or rather ORNNS in general) compute piecewise polynomials, where the semialgebraic decomposition of the domain is similarly defined by the combination of active neurons
$b_{\ell,1}+A_{\ell,1}h_{\ell-1}+B_{\ell,1}(I\otimes X)h_{\ell-1}\geq 0,\qquad\ell=1,\dots,L$
@@
By definition of $ISD(\mathbb{R}^{mn}),$ $f$ is a lattice polynomial $q$ of some $\{p_{1},\dots,p_{J}\}\subset\mathbb{R}[X_{1},\dots,X_{mn}]$. An easy construction shows there exists networks $\mathcal{N}_{1},\dots,\mathcal{N}_{J}$ computing the polynomials $\mathcal{N}_{j}=p_{j}\circ\mathrm{vec}^{-1}$. Another easy construction shows that $MRNN(\mathbb{R}^{m\times n})$ is closed under the lattice operations, so the lattice polynomial $q$ can be computed by a network. In particular, $\mathcal{N}=q\left(\mathcal{N}_{1},\dots,\mathcal{N}_{J}\right)$ is the desired network. ∎
@@
$h_{L}$ $=b_{L,0}+A_{L,0}h_{L-1}+B_{L,0}(I\otimes X)h_{L-1}+$ (25)
$\quad\sigma\left(b_{L,1}+A_{L,1}h_{L-1}+B_{L,1}(I\otimes X)h_{L-1}\right).$
@@
for some $N\in\mathbb{N}$ corresponding to the width of this layer. For every $j=1,\dots,N$, $g_{j}(X)$ is computed by an $L-1$ network, so the inductive hypothesis states there exists $f_{j}\in ISD(\mathbb{R}^{mn})$ such that $g_{j}=f_{j}\circ\mathrm{vec}$. We now rewrite each term of equation (25) using these $f_{j}$. Let $i\in\{0,1\}$.
@@
We can thus rewrite (25) in the equivalent form
@@
Theorems LABEL:thm-MRNN-equals-ISD and LABEL:thm-lattice-over-subring-of-f-ring-is-ring allow us to conclude that $MRNN(\mathbb{R}^{n\times m},\mathbb{R}^{k})$ forms a unital, associative algebra over $\mathbb{R}$; that is, every scalar multiple, sum, and Hadamard product of a MRNNs is computable by an MRNN. Associativity and existence of a multiplicative identity, as well as the ability to compute scalar multiples and sums are obvious and hold for linear ReLU networks as well. What distinguishes MRNNs from linear ReLU networks is the ability to compute Hadamard products. We require this property to build SANNs capable of computing any semialgebraic function.
While the proofs above are not explicitly constructive, they can be made constructive using the identities in Henriksen & Isbell (1962) or Hager & Johnson (2010). Unfortunately, representing the Hadamard product of two MRNNs as an MRNN requires exponentially many parameters, which limits the practical usefulness of a direct application of these constructions in code.
In their original paper introducing the architecture, de Hoop et al. (2022) note that the range of an MRNN is a multivariate piecewise polynomial; we have refined this result to show that MRNNs are ISD. We now ask the natural question, is every piecewise polynomial representable by an MRNN? As noted previously, this question is the famous Pierce–Birkhoff conjecture, a long-standing open problem in real algebraic geometry. MRNNs can represent every piecewise polynomials if and only if the Pierce–Birkhoff conjecture is true.
@@ Appendix C Semialgebraic Functions as Kernels of ISD Functions
[…]
Proposition LABEL:prop-semialgebraic-sets-as-kernels.
@@
Each term $p_{i}^{2}(x)$ (resp. $\max\{0,-q_{i}(x)\}^{2}$) is zero precisely when the corresponding condition $p_{i}(x)=0$ (resp. $q_{i}(x)\geq 0$) is satisfied. Also, the function $q\mapsto\max\{0,q\}^{2}$ is in $C^{1}$ and has the continuous derivative $q\mapsto 2\max\{0,q\}$. Moreover, each term is non-negative, so the sum is non-negative and zero if and only if every term is zero, meaning every condition is satisfied simultaneously. Finally, $f$ is ISD since it is constructed using the ISD lattice-ring operations on polynomials (see Appendix A), and it is $C^{1}$ since each term in the sum is $C^{1}$.
Now let $S$ be an arbitrary closed semialgebraic set; i.e., the union of basic semialgebraic sets. We may take the basic sets to be closed, since
@@
Corollary LABEL:cor-semialgebraic-graphs-as-kernels.
@@ D.1 Semialgebraic Sard’s and Transversality Theorems
In the “textbook” Sard’s theorem, the smoothness requirements on the function depend on the dimension of the range (Abraham & Robbin, 1967). In contrast, for semialgebraic functions continuous differentiability is sufficient for the theorem to hold. A proof can be found in Kurdyka et al. (2000).
@@
A proof can be found in (Abraham & Robbin, 1967), substituting the semialgebraic Sard’s theorem for Smale’s theorem as required.
@@ D.2 Homotopy existence
We now state a standard result in the theory of homotopy continuation for solving nonlinear equations. Theorem LABEL:thm-homotopy-continuation is adapted from Krantz & Parks (2003) with small modifications using the theorems from Appendix D.1 to specialize to $ISD^{1}$ and avoid any $C^{\infty}$ assumptions. It provides sufficient conditions under which we are guaranteed to be able to evaluate a homotopy from $t=0$ until $t=1$.
@@ D.3 Curve tracing
Let $z(s):=\big{(}y(s),t(s)\big{)}$, and $H\in ISD^{1}(\mathbb{R}^{n}\times\mathbb{R},\mathbb{R}^{n})$ satisfy the hypotheses of Theorem LABEL:thm-homotopy-continuation. Then $(y,1)\in\ker H$ can be computed by solving the ODE initial-value problem defined by the “arc-length parameterization” (Chen & Li, 2015):
@@
With proper choice of $\mathcal{N}$ and $c_{\text{max}}$, the curve-tracing ODE system (30)–(33) can be written in the form of the ODE system (6)–(7) that define SANNs. At time $s=0$, write
@@
where $\alpha$ is a constant chosen to satisfy (31) and (32). The $M$ and $b$ defined above appear in the ODE (6)–(7). Thus we have shown for $H$ satisfying the hypotheses of Theorem LABEL:thm-homotopy-continuation, $y_{0}=0$, and large enough $c_{\max},$ there is an $\mathcal{N}\in ISDnet(m,n,k)$ such that the integral curve traced in solving the ODE (6)–(7) is “correct” at time $s=0$; i.e. it is tangent to the integral curve defined by $\dot{z}$ solving the IVP (30)–(33).
For time $s>0$, we are no longer guaranteed that the first $n$ columns of $DH$ form an invertible submatrix. However, by assumption A.3, $DH$ always has rank $n$, so every $s$ belongs to some interval where we can choose a column of $DH$ to be $\widehat{b}$, and the remaining columns to be $\widehat{M}$. The computations in each of these intervals can be glued together in a way consistent with the ODE (6)–(7) used to compute the entire integral curve. The details can be found in the proof of Theorem 12 in Appendix E.
@@ E.1 Notation
The theorems in this section require us to show the existence of ISD $\mathcal{N}$ generating certain initial value problems for (6)–(7). In Algorithm 1, $\dot{z}$ implicitly depends on the inputs $\mathcal{N}$, $c_{\max}$, and $x$, and ODESolve always runs from time $s=0$ to $s=1$. For this section, we make the dependence on all variables explicit by using the signature
@@ E.2 Glueing
For “$t$-glueing,” we build $\mathcal{N}_{3}$ to interpolate between $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ in disjoint closed $t$ intervals $[t_{0},t_{1}]$ and $[t_{2},t_{3}]$. We use $t$-glueing when we can guarantee the integral curve traced by ODESolve will never enter the region $(t_{1},t_{2})$, so the value of $\mathcal{N}_{3}$ may be arbitrary here. We are free to use this interval to interpolate between $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$.
@@
where $u$ is the Hermite interpolant (Süli & Mayers, 2003) such that
@@
Simple calculus verifies conclusion (36), and conclusion (37) holds since $\dot{v}(s)=0$ outside $[s_{1},s_{2}]$. ∎
@@
Choose $\epsilon>0$ such that $s_{0}<s_{1}-\epsilon<s_{1}+\epsilon<s_{\text{final}}$, and apply Lemma 38 to $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ so that both $b_{1}$ and $b_{2}$ are identically 00 in the region $(s_{1}-\epsilon,s_{1}+\epsilon).$ No matter the value of $M_{1}$ and $M_{2}$, both $\dot{y}$ and $\dot{t}$ are identically 00 here. Let $\lambda(s)=(s-(s_{1}-\epsilon))/(2\epsilon)$. Define
@@
$M_{3}$ is clearly continuous, and $b_{3}$ is continuous since $b_{1}$ and $b_{2}$ are both 00 when $s=s_{1}.$ The constructed $\mathcal{N}_{3}$ is ISD and satisfies conclusion (44). ∎
@@
Given $\mathcal{N}_{1}$, $\mathcal{N}_{2}$, $c_{1}$, $c_{2}$, $y_{1}$, $y_{2}$, $t_{1}$, $t_{2}$,$s_{0}$, $s_{\text{final}}$, let
@@
Suppose $\mathcal{T}_{1}\cap\mathcal{T}_{2}=\emptyset$. Then there exists $\mathcal{N}_{3}$, $c_{3}$ such that